The average expenditure per student, based on daily average attendance, for a certain school year was Php 496,176.00, with a population standard deviation of Php 74,880.00.

A survey for the next school year of 150 randomly selected students resulted in a sample average expenditure of Php 518,304.00. Do these results indicate that the average expenditure has decreased? Choose your own level of significance.


1
Expert's answer
2020-10-19T16:29:45-0400

Null hypothesis "H_0: \mu =496176."

Alternative hypothesis "H_a: \mu>496176."

Level of significance: "\alpha =0.05."

Test statistic: "z=\frac{\bar x-mu}{\frac{s}{\sqrt{n}}}=\frac{518304-496176}{\frac{74880}{\sqrt{150}}}=3.62."

P-value: "p=P(Z>3.62)=0.0001."

Since the P-value is less than 0.05, reject the null hypothesis.

There is a sufficient evidence that the average expenditure has increased.